A restaurant has a fixed costs of 137.50 per day and an average unit cost $4.75 for each meal served. If a typical meal cost $ 6 , how many customers must eat at the restuant each day for the owner to break even
Answer by mananth(16946)

Let meals served be x
daily cost = 137.50+4.75x
Revenue= 6x
To break even they have to be equal.
137.50+4.75x=6x
137.50=6x-4.75x
137.50=1.25x
x= 137.50/1.25
x= 110 meals